# SAFE-SQL: Self-Augmented In-Context Learning with Fine-grained Example Selection for Text-to-SQL

### v2

@@ Abstract
Jimin Lee,  Ingeol Baek,  Byeongjeong Kim  Hyunkyung Bae,  Hwanhee Leethanks: Corresponding author. Department of Artificial Intelligence, Chung-Ang University, Seoul, Korea {ljm1690, ingeolbaek, michael97k, hwanheelee}@cau.ac.kr
@@ 4.2 Baselines
SYN-SQL Sense 13B 95.2 88.6 75.9 60.3 83.5 16.34
SQL-Palm PaLM2 93.5 84.8 62.6 48.2 77.3 18.93
@@ 4.6 Analysis
Score cos $\theta$ of Generated EX % Filtered EX
@@
Question1: What are the names, countries, and ages for every singer in descending order of age? SELECT name, country, age FROM singer ORDER BY age DESC What are the names, ages, and countries of all singers from a specific country, sorted by age in descending order? 1. Identify the desired columns: name, age, and country. Semantic similarity = 10 Structural Similarity = 10 Reasoning path = 10 Relevance score = (10+10+10)/3 = 10
Question2: Return the names and template ids for documents that contain the letter w in their description. SELECT document_name, template_id FROM Documents WHERE Document_Description LIKE "%w%" Retrieve the titles and category IDs of articles whose summaries contain the word "data". 1. Identify the necessary columns: extract title and category_id from the Articles table. Semantic similarity = 7 Structural Similarity = 9 Reasoning path = 8 Relevance score = (7+9+8)/3 = 8
Question3: What is the number of car models that are produced by each maker and what is the id and full name of each maker? SELECT Count(*), T2.FullName, T2.id FROM MODEL_LIST AS T1 JOIN CAR_MAKERS AS T2 ON T1.Maker = T2.id GROUP BY T2.id; List all employees who work in the IT department along with their employee ID and hire date. 1. Identify required details: employee ID and hire date. Semantic similarity = 6 Structural Similarity = 3 Reasoning path = 2 Relevance score = (6+3+2)/3 = 3.67
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].

### v3

@@ Abstract
Jimin Lee1,  Ingeol Baek1,  Byeongjeong Kim1,  Hyunkyung Bae2,  Hwanhee Lee1thanks: Corresponding author. 1Department of Artificial Intelligence, Chung-Ang University, 2New York University {ljm1690, ingeolbaek, michael97k, hwanheelee}@cau.ac.kr, hb3135@nyu.edu
@@ 4.2 Baselines
SYN-SQL Sense-13B 95.2 88.6 75.9 60.3 83.5 -
SQL-PaLM PaLM-2 93.5 84.8 62.6 48.2 77.3 -
@@ 4.6 Analysis
Score cos $\theta$ # of Generated EX % Filtered EX
@@
Question1: What are the names, countries, and ages for every singer in descending order of age? SELECT name, country, age FROM singer ORDER BY age DESC What are the names, ages, and countries of all singers from a specific country, sorted by age in descending order?
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)] and a project for Collabo R&D between Industry, University, and Research Institute funded by Korea Ministry of SMEs and Startups in 2025.(RS-2025-02323112).